find vertex of porabola y = 6 x + 7 - x^2
Answer by nerdybill(7384)

y = 6 x + 7 - x^2
rearranging:
y = -x^2 + 6x + 7
.
The x-coordinate of the vertex is at:
x = -b/(2a) = -6/(2*(-1)) = -6/(-2) = 3
.
The y-coordinate:
y = -x^2 + 6x + 7
y = -3^2 + 6(3) + 7
y = -9 + 18 + 7
y = 16
.
Answer: (3, 16)